On a standardized test with a normal distribution the mean is 500 and the deviation is 100. If 1000 people took the test approximately how many would theoretically score between 400 and 600?
Answer by ewatrrr(24785)

 
Hi
Population: mean is 500 and the deviation is 100.
P(400 ≤ x ≤ 600) = P(-1 ≤ z ≤ 1) = P(z ≤ 1) - P(z ≤ -1) = .8413 - .1587 = .6826
For the normal distribution: Below: z = 0, z = ± 1, z= ±2 , z= ±3 are plotted.
Area under the standard normal curve to the left of the particular z is P(z)
Note: z = 0 (x value: the mean) 50% of the area under the curve is to the left and 50% to the right

one standard deviation from the mean accounts for about 68% of the set
two standard deviations from the mean account for about 95%
and three standard deviations from the mean account for about 99.7%.